# Explaining Large Language Models with gSMILE: Generative Statistical Model-Agnostic Interpretability with Local Explanations

### v2

@@ Abstract
Large Language Models (LLMs) such as GPT, LLaMA, and Claude have achieved remarkable performance in text generation, yet their decision-making processes remain opaque. This lack of transparency poses challenges in domains where trust and accountability are critical. In this work, we introduce gSMILE (Generative SMILE), a generalized and LLM-oriented extension of the previously proposed SMILE framework, for interpreting how these models respond to specific parts of an input prompt.
gSMILE is model-agnostic and operates by introducing controlled perturbations to the input, observing the resulting output changes, and identifying which input words have the most significant impact on the output. This process enables the generation of intuitive heatmaps that visually highlight influential words in the prompt. We evaluate gSMILE across several leading large language models (LLMs) using metrics such as accuracy, consistency, stability, and fidelity, demonstrating its ability to provide precise, robust, and reliable explanations. By making LLMs more interpretable, gSMILE takes an important step toward building more transparent and trustworthy AI systems.
Figure 1: Highlighting the contribution of each word in the input prompt of a Commercial LLM
@@ 1 Introduction
Large Language Models (LLMs), such as GPT, Claude, and LLaMA, have significantly advanced the capabilities of natural language processing systems [1, 2, 3]. These models, based on the Transformer architecture [4], are trained on extensive corpora and exhibit remarkable performance across a wide range of tasks, including text generation, summarisation, code completion, and dialogue reasoning [5]. However, despite their widespread adoption, LLMs remain largely opaque in how they make predictions, leading to concerns about transparency, trust, and safety, especially in high-stakes domains such as healthcare and law [6, 7].
The black-box nature of LLMs limits our ability to understand which parts of an input prompt have the most significant influence on the model’s output. This lack of interpretability hinders debugging, ethical assessment, and responsible deployment [8, 9]. Interpretability methods, especially those that operate post hoc and in a model-agnostic fashion, offer a way to open these black boxes by explaining the model’s behavior without requiring access to internal parameters [10].
In this work, we propose gSMILE, a generalized and LLM-oriented extension of the SMILE (Statistical Model-Agnostic Interpretability with Local Explanations) framework [11]. gSMILE interprets model outputs by perturbing input prompts, measuring output shifts using Wasserstein distance, and fitting a local linear surrogate model. The result is a weighted heatmap that visualizes how individual words in a prompt contribute to the model’s response.
We evaluate gSMILE across multiple large language models (LLMs) and prompt types using quantitative metrics, including fidelity, consistency, stability, and accuracy. Our findings demonstrate that gSMILE yields stable, interpretable, and reliable attributions, rendering it suitable for practical applications in real-world scenarios involving instruction-based generation.
@@
  • Visual attribution through weighted heatmaps: The framework generates intuitive heatmaps that highlight high-impact tokens, enabling human-aligned interpretation of model behavior.
  • Robust cross-model evaluation: We evaluate gSMILE on multiple instruction-tuned LLMs using diverse metrics, accuracy, fidelity, stability, and consistency that demonstrate its reliability and adaptability.
@@ 2.1 Large Language Models
Large language models (LLMs) have become increasingly prevalent due to significant advancements in deep learning [12], improved access to large-scale datasets, and the growing availability of powerful computing resources [13]. These models are typically trained on massive corpora of text. They are capable of generating fluent, contextually appropriate responses across a wide variety of language-based tasks, including question answering, summarisation, translation, and content generation. Over time, several influential large language models (LLMs) have helped push the field forward, each introducing novel architectural innovations [4], training strategies [14], or alignment techniques [15] that have contributed to their growing capabilities and broader adoption.
Among the most influential models, GPT (OpenAI) [1] stands out as a leading autoregressive transformer-based architecture capable of generating coherent and contextually rich text. It set a new standard for generative language models and introduced the broader public to powerful applications of AI through tools like ChatGPT. Expanding on this paradigm, Google’s T5 [14] reframed all NLP tasks as text-to-text problems, streamlining the transfer of learning across diverse tasks and contributing to more unified and efficient architectures. Palm (Google) [16], with its Pathways design, extended the multitask learning capability at scale, allowing a single model to generalize effectively across different NLP applications.
In parallel, Anthropic’s Claude [17] introduced a new approach centered on safe and aligned AI behavior. Claude utilizes constitutional AI principles to ensure its outputs are consistent with predefined ethical standards, thereby promoting greater transparency and reducing the risk of harm. More recently, Google’s Gemma 2 [18] has focused on logical reasoning and structured data interpretation, offering a competitive option for research-driven and enterprise-level deployment. Meta’s LLaMA [19] series continues to draw attention as a scalable, open-source alternative, providing the research community with accessible tools that deliver high performance across a range of benchmarks.
Community-driven efforts have also enriched the landscape. EleutherAI’s GPT-NeoX [20] was developed to provide an open-source alternative to proprietary GPT models, offering researchers greater flexibility in experimentation and deployment. AI21 Labs’ Jurassic-2 [21] targets commercial-scale applications with a focus on contextual awareness and enterprise integration. Meanwhile, Cohere’s Command R+ [22] integrates retrieval-augmented generation (RAG), enhancing factual consistency and minimizing hallucinations in the generative output — a key concern for the practical adoption of large language models (LLMs).
Additionally, transformer-based models, such as BERT [23], have revolutionized NLP through bidirectional encoding, providing substantial improvements in tasks like question answering, named entity recognition, and sentiment analysis. Efficiency-focused models, such as Microsoft’s Phi-4, highlight a growing trend toward parameter-efficient design without sacrificing core performance, making large language models (LLMs) more accessible for low-resource environments. Mistral and Qwen 2.5 (by Alibaba) continue this direction by offering fine-tuning-friendly models with strong multilingual and domain adaptation capabilities.
Several recent open-source contenders further enrich the ecosystem. DeepSeek-LLM [24] delivers performance comparable to proprietary models while remaining open and extensible. Mixtral and StableLM focus on task efficiency and targeted performance, ensuring their utility in production environments. DeepSeek-R1 has been noted for its emphasis on reasoning, reflecting a broader shift from general language generation to models optimized for critical thinking, chain-of-thought prompting, and interpretability.
Cutting-edge advancements have introduced even more specialized architectures. Qwen2.5-Max utilizes large-scale Mixture-of-Experts (MoE) models to enhance throughput and scalability. OpenAI’s o3-mini emphasizes low-latency, cost-effective deployment while retaining high-quality output — ideal for mobile and edge computing. Perhaps most notably, DeepSeek-V3 represents a significant milestone by achieving GPT-4o-level performance in an open-source setting, suggesting that open-source models may soon rival or surpass closed models in both capability and adoption.
The following models in Fig. 2 are explicitly designed for text-to-text tasks, meaning they take text input and generate text output across various NLP applications. These models continue to define the state of the art in machine-generated language, with a growing focus on transparency, efficiency, and alignment with human values.
@@
Category Models Description General-Purpose GPT-4 [25], GPT-4o [26], Claude 3 [27], Gemini 1.5 [28], Command R+ [29], Mistral [30] Versatile for wide range of NLP tasks Open-Source LLaMA 3 [31], Mistral [30], Mixtral [32], Falcon 180B [33], DeepSeek-V3 [34], OpenChat [35] Customizable models for research applications Retrieval-Augmented Command R+ [29], GPT-4 Bing [25], Claude 3 RAG [27], Gemini Search [28] Enhance information retrieval accuracy, reduce misinformation Long-Text Processing Claude 3 [27], Gemini 1.5 [28], GPT-4o [26], Grok-1.5 [36] Specialized in processing long-form text Advanced Reasoning GPT-4 [25], Claude 3 Opus [27], Gemini 1.5 [28], Mixtral [32], Grok-1.5 [36] Logical reasoning, multi-task learning, problem-solving Safety-Focused Design Claude 3 [27], Gemini [28], GPT-4 [25], Grok [36] Emphasize alignment with safety rules and bias mitigation Mixture-of-Experts (MoE) Mixtral [32], Qwen2.5-Max [37], GShard [38], Switch Transformer [39] Improved efficiency and performance via MoE Multimodal GPT-4o [26], Gemini 1.5 [28], Claude 3 [27], Kosmos-2 [40], Grok-1.5V [36] Capable of processing both text and visual inputs Multilingual BLOOM [41], XGLM [42], mBERT [43], XLM-R [44], ByT5 [45] Optimized for cross-lingual understanding and translation Domain-Specific Med-PaLM 2 [46], Galactica [47], Legal-BERT [48], StarCoder2 [49], Code LLaMA [50] Tailored for specific tasks such as medical, legal, or code Instruction-Tuned Alpaca [51], Vicuna [52], ChatGLM [53], Baize [54], OpenChat [35] Tuned for better instruction-following and chat capabilities Lightweight / Edge Phi-2 [55], TinyLLaMA [56], DistilGPT2 [57], GPT-2 [58], LLaMA 2 7B [59] Designed for fast inference and deployment on edge devices
@@ 2.2 Explainable AI (XAI)
Explainability methods can be broadly categorized as intrinsic or post-hoc. Intrinsic methods involve inherently interpretable models due to their simplicity, such as linear regression or decision trees [62]. These models are often preferred in applications where transparency is critical, but they may lack the predictive power of more complex models [6].
@@
A key improvement in LIME’s explanations is the refinement of distance measures used to weigh data points. For instance, SMILE uses statistical techniques to generate more consistent and reliable explanations. However, while these improvements enhance reliability, they also come with increased computational complexity. Another approach to improving LIME focuses on upgrading its simple linear models with more sophisticated alternatives. Methods such as Q-LIME [67], S-LIME [68], Bay-LIME, and ALIME use more advanced surrogate models to provide explanations that better reflect the underlying model’s behavior, capturing non-linear relationships that the original method might miss. Furthermore, researchers have explored optimization strategies to enhance the efficiency and stability of LIME. Techniques like OptiLIME and G-LIME work to strike a balance between accuracy and computational cost, ensuring that explanations remain reliable while keeping the process computationally manageable.
@@
Recent research has extended LIME to support specific data types and improve its interpretability. Adaptations such as PointNet LIME for 3D point cloud data [69], TS-MULE for time series data [70], Graph LIME for graph-structured data [71], Sound LIME for audio data [72], and B-LIME for ECG signal data [73] demonstrate how LIME can be tailored to diverse applications. In addition, several modifications have been proposed to enhance LIME’s approach, including the use of alternative surrogate models, fine-tuning distance parameters, optimizing sampling techniques, and refining computational efficiency to improve both interpretability and accuracy.
@@
Other extensions include DSEG-LIME [74], which tailors LIME for segmentation models by generating perturbations at the object level rather than the pixel level, improving interpretability in image segmentation tasks. SLICE [75] introduces structured sampling and clustering to ensure diverse and representative local explanations, reducing redundancy in LIME samples. Stratified LIME [76] improves fidelity by stratifying the input space and generating more targeted local samples for different subpopulations. US-LIME [77] employs uncertainty sampling to guide instance selection, increasing the reliability and faithfulness of local explanations. SS-LIME [78] uses self-supervised learning to identify important regions in the feature space and provide richer explanations. K-LIME [79] incorporates kernel-based methods into LIME’s regression step to model more complex local decision boundaries. Attention-LIME [80] leverages attention mechanisms to prioritize features during the explanation, aligning more closely with deep learning models. BMB-LIME [81] employs a boosted multi-branch approach to combine multiple local models, thereby enhancing consistency and robustness.
@@
These methods aim to enhance interpretability across various domains, including segmentation, survival analysis, multimodal learning, and attention-based contexts. Additionally, a wide range of other extensions have been developed to enhance LIME’s capabilities across domains and explanation strategies, including: LSLIME, LIME-Aleph, Anchors [82], LIME-SUP, LORE, ILIME, Aleph LIME, bLIMEy, NormLIME, DLIME, LIMEtree, MPS-LIME, Explain Explore, survLIME, survLIME-inf, CBR-LIME, GuidedLIME, audioLIME, MeLIME, C-LIME, gLIME, DIME, s-LIME, LIMESegment, UnRAvEL-LIME, Kernel-LIME, Specific-Input LIME, ExpLIMEable, GLIME, GMM-LIME, SEGAL, and Sig-LIME, reflecting the diverse and rapidly evolving landscape of post-hoc explainability [83].
@@ 2.3 Explainable AI (XAI) for Large Language Models
After training, post-hoc explainability methods aim to interpret LLMs, providing insights into their decision-making processes [84, 86]. Zhao et al. [84] present a taxonomy of explainability techniques for LLMs, categorising approaches into local and global explanations, and discussing limitations and future research directions. Similarly, Bender et al. [85] highlight the risks of opaque LLMs, including biases and misinformation, and call for increased transparency and accountability. Recent surveys further expand this field by reviewing how LLMs contribute to explainable AI at large [87, 88]. Amara et al. [89] recently introduced ConceptX, a concept-level attribution method highlighting semantically meaningful input phrases through a coalition-based Shapley framework, offering robust and human-aligned explanations.
@@
Mechanistic interpretability seeks to uncover how LLMs process information internally [90, 91, 92]. Olah et al. [90] propose reverse-engineering techniques to dissect internal representations. Nanda et al. [91] examine how LLMs develop abstract representations and generalisation, introducing progress measures for interpretability research. Recent work by Anthropic [92] demonstrates the role of mechanistic interpretability in understanding and controlling model behaviour. Additionally, the SEER framework [93] introduces self-explainability mechanisms to enhance the interpretability of internal representations in LLMs. The CELL framework [94] further advances mechanistic interpretability by integrating concept-based explanations directly into the training and representation learning of language models.
@@
Recent advancements from the Transformer Circuits framework have further enriched mechanistic interpretability. Using causal scrubbing, Meng et al. [95] introduced Attribution Graphs to uncover emergent structures within neural networks. In contrast, Nanda et al. [96] investigated the scaling of monosemanticity, extracting interpretable features from state-of-the-art models. Additional recent contributions explore specific tasks such as relevance estimation [97], emotion inference [98], and geospatial reasoning [99]. These works leverage techniques like activation patching and sparse autoencoders to reveal LLM internals. Sparse interpretability methods are further reviewed comprehensively in [100]. Neuroscientific perspectives have also been applied, using dynamical systems theory to model token-level trajectories within transformers [101].
@@
Human-centric explainability focuses on making LLM outputs understandable to non-experts [102, 103, 104]. Ji et al. [102] investigate hallucination in LLM outputs and propose techniques for detecting and mitigating incorrect responses. Krause et al. [103] assess reasoning abilities compared to human performance, and Martens et al. [104] explore counterfactual explanations to improve human understanding. Recent works extend this perspective by enabling interactive explanations of black-box models through fine-tuned LLMs [105] and by benchmarking explanation quality in clinical domains [106]. These approaches improve the accessibility and transparency of LLM outputs in real-world decision-making.
@@
Explainability enhances debugging, bias detection, regulatory compliance, and user trust in LLMs [107]. Challenges remain, including defining “meaning” in LLM-generated outputs [107], managing the ethical and societal implications of black-box models, and balancing performance with interpretability. Recent work has demonstrated SMILE’s applicability beyond language, such as in instruction-based image editing [108], highlighting its modality-agnostic nature. New efforts address rare concept discovery via tailored sparse autoencoders [109] and propose hybrid architectures like concept bottleneck models [110] and symbolic program compositions [111] to enable more transparent reasoning. Future research should emphasise standardised benchmarks, hybrid mechanistic and human-centric methods, and integrating explainability into AI governance frameworks.

### v4

@@ Abstract
Large Language Models (LLMs) such as GPT, LLaMA, and Claude achieve remarkable performance in text generation but remain opaque in their decision-making processes, limiting trust and accountability in high-stakes applications. We present gSMILE (generative SMILE), a model-agnostic, perturbation-based framework for token-level interpretability in LLMs. Extending the SMILE methodology, gSMILE uses controlled prompt perturbations, Wasserstein distance metrics, and weighted linear surrogates to identify input tokens with the most significant impact on the output. This process enables the generation of intuitive heatmaps that visually highlight influential tokens and reasoning paths. We evaluate gSMILE across leading LLMs (OpenAI’s gpt-3.5-turbo-instruct, Meta’s LLaMA 3.1 Instruct Turbo, and Anthropic’s Claude 2.1) using attribution fidelity, attribution consistency, attribution stability, attribution faithfulness, and attribution accuracy as metrics. Results show that gSMILE delivers reliable human-aligned attributions, with Claude 2.1 excelling in attention fidelity and GPT-3.5 achieving the highest output consistency. These findings demonstrate gSMILE’s ability to balance model performance and interpretability, enabling more transparent and trustworthy AI systems.
Figure 1: Overview of gSMILE interpretability process: An input prompt is processed by a transformer-based LLM, and gSMILE generates a token-level heatmap highlighting influential words contributing to the output.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 [1], Claude [2], and LLaMA [3] have rapidly become central to artificial intelligence, powering applications in healthcare, education, law, and scientific research [4, 5, 6, 7]. Built on the Transformer architecture [8] and trained on massive corpora, they generate fluent, context-aware responses for tasks ranging from dialogue and summarisation to reasoning and information extraction [6].
Today, LLMs are embedded in widely used technologies, assisting clinicians in healthcare [4], supporting personalised education [5], streamlining customer service [6], and aiding legal research [7]. Their growing influence underscores both their transformative potential and the urgent need to address concerns about reliability, transparency, and trust.
Despite their remarkable capabilities, a fundamental question remains: how much control do we truly have over LLM behaviour, and to what extent do we understand their decision-making processes? While they can generate compelling outputs, studies show that these responses are not always reliable, often containing hallucinations or logical inconsistencies that damage trust [4, 5]. In practice, users are left uncertain about why a model provides one answer instead of another, raising concerns about whether we genuinely comprehend the mechanisms driving their outputs [9].
Recent work has also highlighted the critical role of prompt design in shaping LLM performance. Simple changes to instructions, for example, adding phrases such as “Let’s think step by step” or “Take a deep breath and work on this problem step-by-step, have been shown to improve accuracy on reasoning benchmarks significantly. OPRO reports that the latter phrase boosts GSM8K accuracy to 80.2%, compared to 71.8% with the baseline step-by-step prompt [10]. Similarly, Self-Supervised Prompt Optimisation (SPO) demonstrates that optimised prompts, even without external references, consistently improve results across multiple datasets [11]. These findings underline the power of prompt engineering: if we understood more clearly which features of a prompt drive performance, we could design instructions that yield more reliable outputs. Interpretability methods can provide the insights needed to guide such prompt design.
Moreover, interpretability approaches must be model-agnostic. While Anthropic has recently showcased advanced interpretability research for Claude 3.5, using techniques such as circuit tracing and attribution graphs to visualise how the model plans, manipulates concepts, and sometimes even fabricates reasoning [12], such access to internal mechanisms is limited to others with privileged model access. For most practitioners and end users, leading LLMs like GPT and Claude are available only through APIs that restrict visibility into their parameters and architectures. As a result, practical interpretability techniques must be designed to explain model behaviour without relying on direct access to internal weights or structural details [6].
To this end, we propose gSMILE, a generalised and LLM-oriented extension of the SMILE (Statistical Model-Agnostic Interpretability with Local Explanations) framework [13]. gSMILE interprets model outputs by perturbing input prompts, measuring output shifts using Wasserstein distance, and fitting a local linear surrogate model. The result is a weighted heatmap that visualises how individual words in a prompt contribute to the model’s response.
We evaluate gSMILE across multiple large language models (LLMs) and prompt types using quantitative metrics, including ATT fidelity, ATT consistency, ATT stability, ATT faithfulness and ATT accuracy. Our findings demonstrate that gSMILE yields stable, interpretable, and reliable attributions, rendering it suitable for practical applications in real-world scenarios involving instruction-based generation.
@@
  • Visual attribution through weighted heatmaps: The framework generates intuitive heatmaps that highlight high-impact tokens, enabling human-aligned interpretation of model behaviour.
  • Robust cross-model evaluation:We evaluate gSMILE on multiple instruction-tuned LLMs using diverse metrics, including ATT accuracy, ATT fidelity, ATT stability, ATT faithfulness and ATT consistency, that demonstrate its reliability and adaptability.
List of Abbreviations
Abbreviation Definition
LLM Large Language Model
XAI Explainable Artificial Intelligence
SMILE Statistical Model-agnostic Interpretability with Local Explanations
gSMILE Generative SMILE (proposed method)
WMD Word Mover’s Distance
IWMD Inner Word Mover’s Distance (input-level distance)
OWMD Outcome Word Mover’s Distance (output-level distance)
WD Wasserstein Distance
ATT Attribution
TV Total Variation distance
P Norm order in Wasserstein-$P$ metric
$\pi^{(n)}$ Black-box text generation model
$\Delta(x,\hat{x}_{j})$ Output-level semantic shift
$w_{j}$ Similarity-based kernel weight
$z_{j}$ Feature vector representation of perturbed prompt
@@ 2.1 Large Language Models
Large language models (LLMs) have rapidly advanced due to progress in deep learning [14], the availability of massive datasets [15], and increasingly powerful hardware. Trained on large text corpora, they can generate fluent responses for tasks such as question answering, summarisation, translation, and content creation. Key innovations include new architectures [8], training strategies [16], and alignment methods [17].
Several models have shaped the field. OpenAI’s GPT series [18] popularised autoregressive transformers through ChatGPT. Google’s T5 [16] reframed NLP as a text-to-text problem, while PaLM [19] scaled multitask learning. Anthropic’s Claude [12] introduced constitutional AI to emphasise safety, and Google’s Gemma [20] advanced reasoning-focused applications. Meta’s LLaMA models [3] provided a scalable open-source alternative.
Community-driven projects also played a significant role. EleutherAI’s GPT-NeoX [21] offered an open substitute for proprietary GPT models. AI21’s Jurassic-2 [22] targeted enterprise integration, and Cohere’s Command R+ [23] combined generation with retrieval. Earlier, BERT [24] transformed bidirectional understanding. More recent lightweight models such as Phi-4, Mistral, and Alibaba’s Qwen 2.5 focus on efficiency, multilinguality, and adaptability.
The latest generation continues this trend. DeepSeek-LLM [25], Mixtral [26], and StableLM prioritise open and efficient deployment, while DeepSeek-R1 highlights reasoning. Others push specialisation: Qwen2.5-Max uses Mixture-of-Experts for scalability, OpenAI’s o3-mini is designed for low-latency use, and DeepSeek-V3 has reached GPT-4-level performance in an open-source setting.
As shown in Fig. 2, today’s LLMs are not only powerful text generators but also increasingly efficient, transparent, and aligned with human values. They continue to shape the state of the art in both research and industry.
@@
Category Models Description General-Purpose GPT-4 [1], GPT-4o [27], Claude 3 [28], Claude 2.1 [29] Gemini 1.5 [35], Command R+ [23], Mistral [30] Versatile for wide range of NLP tasks Open-Source LLaMA 3 [31], Mistral [30], Mixtral [26], Falcon 180B [32], DeepSeek-V3 [33], OpenChat [34] Customizable models for research applications Retrieval-Augmented Command R+ [23], GPT-4 Bing [1], Claude 3 RAG [28], Gemini Search [35] Enhance information retrieval accuracy, reduce misinformation Long-Text Processing Claude 3 [28], Gemini 1.5 [35], GPT-4o [27], Grok-1.5 [36] Specialized in processing long-form text Advanced Reasoning GPT-4 [1], Claude 3 Opus [28], Gemini 1.5 [35], Mixtral [26], Grok-1.5 [36] Logical reasoning, multi-task learning, problem-solving Safety-Focused Design Claude 3 [28], Gemini [35], GPT-4 [1], Grok [36] Emphasize alignment with safety rules and bias mitigation Mixture-of-Experts (MoE) Mixtral [26], Qwen2.5-Max [37], GShard [38], Switch Transformer [39] Improved efficiency and performance via MoE Multimodal GPT-4o [27], Gemini 1.5 [35], Claude 3 [28], Kosmos-2 [40], Grok-1.5V [36] Capable of processing both text and visual inputs Multilingual BLOOM [41], XGLM [42], mBERT [24], XLM-R [43], ByT5 [44] Optimized for cross-lingual understanding and translation Domain-Specific Med-PaLM 2 [45], Galactica [46], Legal-BERT [47], StarCoder2 [48], Code LLaMA [49] Tailored for specific tasks such as medical, legal, or code Instruction-Tuned Alpaca [50], Vicuna [51], ChatGLM [52], Baize [53], OpenChat [34], GPT-3.5 [54], LLaMA3.1 [55] Tuned for better instruction-following and chat capabilities Lightweight / Edge Phi-2 [56], TinyLLaMA [57], DistilGPT2 [58], GPT-2 [59], LLaMA 2 7B [3] Designed for fast inference and deployment on edge devices
@@ 2.2 Explainable AI (XAI)
Explainability methods can be broadly categorised as intrinsic or post-hoc. Intrinsic methods involve inherently interpretable models due to their simplicity, such as linear regression or decision trees [62]. These models are often preferred in applications where transparency is critical, but they may lack the predictive power of more complex models [63].
@@
A key improvement in LIME’s explanations is the refinement of distance measures used to weigh data points. For instance, SMILE uses statistical techniques to generate more consistent and reliable explanations. However, while these improvements enhance reliability, they also come with increased computational complexity. Another approach to improving LIME focuses on upgrading its simple linear models with more sophisticated alternatives. Methods such as Q-LIME [68], S-LIME [69], Bay-LIME [70], and ALIME [71] use more advanced surrogate models to provide explanations that better reflect the underlying model’s behaviour, capturing non-linear relationships that the original method might miss. Furthermore, researchers have explored optimisation strategies to enhance the efficiency and stability of LIME. Techniques like OptiLIME [72] and G-LIME [73] work to strike a balance between accuracy and computational cost, ensuring that explanations remain reliable while keeping the process computationally manageable.
@@
Recent research has extended LIME to support specific data types and improve its interpretability. Adaptations such as PointNet LIME for 3D point cloud data [74], TS-MULE for time series data [75], Graph LIME for graph-structured data [76], Sound LIME for audio data [77], and B-LIME for ECG signal data [78] demonstrate how LIME can be tailored to diverse applications. In addition, several modifications have been proposed to enhance LIME’s approach, including the use of alternative surrogate models, fine-tuning distance parameters, optimising sampling techniques, and refining computational efficiency to improve both interpretability and accuracy.
@@
Other extensions include DSEG-LIME [79], which tailors LIME for segmentation models by generating perturbations at the object level rather than the pixel level, improving interpretability in image segmentation tasks. SLICE [80] introduces structured sampling and clustering to ensure diverse and representative local explanations, reducing redundancy in LIME samples. Stratified LIME [81] improves fidelity by stratifying the input space and generating more targeted local samples for different subpopulations. US-LIME [82] employs uncertainty sampling to guide instance selection, increasing the reliability and faithfulness of local explanations. S-LIME [83] uses self-supervised learning to identify important regions in the feature space and provide richer explanations.
@@
These methods aim to enhance interpretability across various domains, including segmentation, survival analysis, multimodal learning, and attention-based contexts. Additionally, a wide range of other extensions have been developed to enhance LIME’s capabilities across domains and explanation strategies, including: LSLIME, LIME-Aleph, Anchors [84]. Numerous other LIME-based and post-hoc explanation variants have been proposed, encompassing a wide range of methodological innovations. These have been comprehensively reviewed and categorised in [85].
@@ 2.3 Explainable AI (XAI) for Large Language Models
After training, post-hoc explainability methods aim to interpret LLMs, providing insights into their decision-making processes [86, 88]. Zhao et al. [86] present a taxonomy of explainability techniques for LLMs, categorising approaches into local and global explanations, and discussing limitations and future research directions. Similarly, Bender et al. [87] highlight the risks of opaque LLMs, including biases and misinformation, and call for increased transparency and accountability. Recent surveys further expand this field by reviewing how LLMs contribute to explainable AI at large [86]. Amara et al. [89] recently introduced ConceptX, a concept-level attribution method highlighting semantically meaningful input phrases through a coalition-based Shapley framework, offering robust and human-aligned explanations.
@@
Mechanistic interpretability seeks to uncover how LLMs process information internally [90, 28]. Olah et al. [90] propose reverse-engineering techniques to dissect internal representations. Nanda et al. [90] examine how LLMs develop abstract representations and generalisation, introducing progress measures for interpretability research. Recent work by Anthropic [28] demonstrates the role of mechanistic interpretability in understanding and controlling model behaviour. Additionally, the SEER framework [91] introduces self-explainability mechanisms to enhance the interpretability of internal representations in LLMs. The CELL framework [92] further advances mechanistic interpretability by integrating concept-based explanations directly into the training and representation learning of language models.
@@
Recent advancements from the Transformer Circuits framework have further enriched mechanistic interpretability. Using causal scrubbing, Ameisen et al. [93] introduced Attribution Graphs to uncover emergent structures within neural networks. In contrast, Nanda et al. [94] investigated the scaling of monosemanticity, extracting interpretable features from state-of-the-art models. Additional recent contributions explore specific tasks such as relevance estimation [95], emotion inference [96], and geospatial reasoning [97]. These works leverage techniques like activation patching and sparse autoencoders to reveal LLM internals. Sparse interpretability methods are further reviewed comprehensively in [95]. Neuroscientific perspectives have also been applied, using dynamical systems theory to model token-level trajectories within transformers [98].
@@
Human-centric explainability focuses on making LLM outputs understandable to non-experts [99, 100, 101]. Ji et al. [99] investigate hallucination in LLM outputs and propose techniques for detecting and mitigating incorrect responses. Krause et al. [100] assess reasoning abilities compared to human performance, and Martens et al. [101] explore counterfactual explanations to improve human understanding. Recent works extend this perspective by enabling interactive explanations of black-box models [101] and by benchmarking explanation quality in clinical domains [102]. These approaches improve the accessibility and transparency of LLM outputs in real-world decision-making.
Recent research has explored cost-efficient, model-agnostic explanation methods for LLMs [103], which align well with the objectives of gSMILE; their approach leverages sampling and scoring strategies to generate faithful explanations without requiring access to model internals.
@@
Explainability enhances debugging, bias detection, regulatory compliance, and user trust in LLMs [60]. Challenges remain, including defining “meaning” in LLM-generated outputs [60], managing the ethical and societal implications of black-box models, and balancing performance with interpretability. Recent work has demonstrated SMILE’s applicability beyond language, such as in instruction-based image editing [104], highlighting its modality-agnostic nature.